OpenStudy (anonymous):

x+5y=4 please graph and show work.

OpenStudy (mathteacher1729):

Step 1: let x = 0, solve for y. Step 2: plot (0, y) Step 3: let y = 0 , solve for x. Step 4: plot (x,0) Step 5: Draw a straight line between the points in step (2) and step (4).

OpenStudy (anonymous):

To find the x intercept let y=0 and solve for x x + 5(0) = 4 x = 4 So the x intercept is 4. Then find the y intercept by letting x=0 0 + 5y = 4 5y / 5 = 4/5 y = 4/5 Now you can pick a number for x to find a third coordinate. Let x=2 x + 5y = 4 2 + 5y = 4 2 - 2 + 5y = 4 - 2 5y = 2 5y / 5 = 2 /5 y = 2/5 Now you have 3 coordinates you can plot. Those coordinates are (4, 0) , (0, 4/5) and (2, 2/5)

OpenStudy (anonymous):

Or you could just simply choose 3 different x coordinates, substitute those numbers into the equation and solve for y and it will give you 3 coordinates to plot. For example: x = 1 x = -6 x = 7 Solve for y by substituting the values of x into the equation as follows: x + 5y = 4 (1) + 5y = 4 1 -1 + 5y = 4 -1 5y = 3 5y/5 = 3/5 y = 3/5 or in decimal form (0.6) x + 5y = 4 (-6) + 5y = 4 (-6) + 6 + 5y = 4 + 6 5y = 10 5y/5 = 10/5 y = 2 x + 5y = 4 7 + 5y = 4 7-7 + 5y = 4 - 7 5y = -3 5y/5 = -3/5 y = -3/5 or in decimal form (0.6) Now you have 3 coordinate sets: (1, 3/5) (-6, 2) (7, -3/5)

OpenStudy (anonymous):

Either way, you get the same line. Hope this helps